# Advancing Multi-Organ Disease Care: A Hierarchical Multi-Agent Reinforcement Learning Framework

### v1

@@ Introduction
In addition to the agent options discussed above for higher-level agents, this section provides detailed information on the actions of leaf-level agents. According to our hierarchical task decomposition approach, each agent operates within a factored action space. It helps reduce the learning burden on each agent and prevents the effects of potentially low-quality samples across subspaces. Based on clinical expertise, we selected seven treatments. Two treatments were considered for the Cardiovascular system: IV fluids (IV) and vasopressors (Vaso), three for Neuro: anesthetics (S1), analgesics (S2), and sedatives (S3), and two for Renal: diuretics and dialysis. Proposed hierarchy is generic which allows integration of agents operating on both continuous and discrete action spaces. However, we tested the approach using discrete action spaces. Dosages for each treatment were discretized into five levels (no-action + 4 quantiles), except dialysis, which is a binary action (active or inactive). Proposed hierarchy supports mixing of any treatment within or across systems, aligning with common clinical practices derived from our data analysis and clinical consultations. For example, we enforced exclusive use of one renal treatment at a time—either diuretics or dialysis, and allowed all other treatment combinations. This showcases the hierarchy’s adaptability to incorporate such domain specific constrains. Accordingly, our solution flexibly handles organ systems with different degrees of complexity and configurations of actions, enabling the effectiveness of the solution in real, complex clinical settings.
RL Reward:
We used a hybrid reward system, combining a mortality-based terminal reward—positive ($+R$) for survival, and negative ($-R$) for death—with clinically guided intermittent rewards to adjust rewards based on immediate health outcomes. This approach addresses reward sparsity in long sequences (see Appendix, Section RL Reward).
Q Learning
@@
Traditional RL frameworks modeled by Markov Decision Processes (MDPs), often assume that actions occur instantaneously with uniform execution times. This assumption simplifies modeling, but fails to capture the complexities of real-world scenarios with actions that span multiple steps and vary in duration.
@@
where $\gamma$ is a discount factor and $\epsilon^{\pi}(o,s,t)$ is the execution of option $o$ in state $s$ from time $t$ to its termination. Unlike standard options framework, our proposed solution contains both independent and cooperative agents controlling options, as detailed in Hierarchical Decomposition section. Thus, the root level Q-values are updated as follows:
$Q(s,o)\leftarrow Q(s,o)+\alpha\left[Q_{tgt}(s,o)-Q(s,o)\right]$ (1)
@@
$Q_{tgt}(s,o)=$$\begin{cases}r,&\text{if }o=A_{Rt}^{0}\\ \max\limits_{a\in A_{Neu}}Q_{Neu}(s,a),&\text{if }o=A_{Rt}^{Neu}\\ \max\limits_{a\in A_{Car}}Q_{Car}(s,a),&\text{if }o=A_{Rt}^{Car}\\ \max\limits_{a\in A_{Ren}}Q_{Ren}(s,a),&\text{if }o=A_{Rt}^{Ren}\\ \max\limits_{a\in A_{OMix}}Q_{OMix}(s,a),&\text{if }o=A_{Rt}^{OMix}\end{cases}$
where $r$ is the immediate reward received for choosing no-action ($A_{Rt}^{0}$), and $A_{Neu}$, $A_{Car}$, and $A_{Ren}$ are the action spaces for the Neuro, Cardio and Renal agents, whereas $A_{Rt}^{OMix}$ is their combined action space ($A_{Neu}\times A_{Car}\times A_{Ren}$). Value functions $Q_{Neu}(s,a)$, $Q_{Car}(s,a)$, and $Q_{Ren}(s,a)$ for the corresponding agents are updated via Temporal Difference (TD) learning as follows:
$Q_{Neu}(s,a)\leftarrow$ $Q_{Neu}(s,a)+\alpha\left[y^{Neu}-Q_{Neu}(s,a)\right]$ (2)
@@
$M_{OMix}$ is trained using the QMix architecture (Rashid et al. 2020), by combining the individual value functions of its sub-agents ($M_{Omix}^{Neu}$,$M_{Omix}^{Car}$, and $M_{Omix}^{Ren}$) into a unified value function ($Q_{OMix}$). QMix framework supports cooperative training and consistent decision-making across agents by enforcing a monotonicity constraint. The constraint keeps the weights of the mixing network non-negative, ensuring that the $argmax$ operation on $Q_{OMix}$ is consistent with those from the sub-agents. Hence, $Q_{OMix}$ is represented as follows:
$\mathrm{arg\,max}_{a\in A_{OMix}}Q_{OMix}(s,a)=$ (3)
[…]
@@
Training process:
We used a two phase approach to train the proposed hierarchical model. In phase one, the root agent ($M_{Rt}$) is trained first. Resulting latent embeddings $E^{Rt}$ are then fine-tuned during the subsequent training of organ-specific agents ($M_{Neu}$, $M_{Car}$, and $M_{Ren}$). In the process, targeted embeddings ($E^{Neu}$, $E^{Car}$ and $E^{Ren}$) are learned and formulate Targeted State Representations. These representations are used to train the corresponding lower-level sub-agents. Independent agents directly learn $Q$ values from received rewards (see Equation 2),whereas cooperative agents $M_{Omix}^{Neu}$,$M_{Omix}^{Car}$, $M_{Omix}^{Ren}$ are trained using shared rewards (Equation 3). In phase two, we integrate the full hierarchy by using the trained organ-specific agents from phase one. $Q(s,a)$ values from these agents are used to retrain $M_{Rt}$ (see Equation 1), and no other agent is retrained. After training, at each timestep, $M_{Rt}$ evaluates the patient state $s_{t}^{Rt}$ and select either no-action, or invokes lower level agents for final treatment and dosage recommendations.
Experiments
Dataset
We collected data on 30,440 patients under Sepsis-3 criteria from the popular MIMIC-IV database (Johnson et al. 2020). Data spans from 24 hours pre-diagnosis to 48 hours post-diagnosis forming a maximum 72 hour window per patient, barring death or ICU discharge. Patient state data collected includes 48 physiological measurements including vital signs, laboratory test results, severity scores and demographics (see Appendix, Section Dataset). State and actions data were aggregated into four-hourly windows to generate uniform patient data sequences. The patient trajectories were randomly split into 75% training and 25% testing sets.
Baselines
We evaluate our model performance against a variety of baselines, including both single- (D3QN-S and SoftAC-S) and multi-agent systems under independent (D3QN-O) and cooperative (QMix-O and QMix-T) learning approaches:
@@
  • Single D3QN (D3QN-S) (Raghu et al. 2017): State-of-the-art single-agent Dueling DQN predicting all action combinations in a flattened action space. We used the original implementation and tuned hyperparameters, and simplified the problem to only 3 quantiles per treatment.
  • Single SoftAC (SoftAC-S) (Haarnoja et al. 2018): A single-agent Soft Actor-Critic predicting all action combinations in a flattened action space. We used the original implementation and tuned hyperparameters, and simplified the problem to only 3 quantiles per treatment.
  • Organ-specific D3QN (D3QN-O): Three independent D3QN agents for Neuro, Cardio and Renal systems. Models are trained using all available samples, including those treated exclusively for the target organ and those mixed with other organ treatments. We present average quantitative metric values across agents for comparisons.
  • Treatment-specific D3QN (D3QN-T): Six independent D3QN agents for S1, S2, S3, IV, vaso, and diuretics/dialysis. Models are trained using all available samples, including those using single-treatments and mixed-treatments. Average quantitative metrics are reported.
@@
  • Treatment-coordinated QMix (QMix-T): Uses six cooperative treatment-level agents (S1, S2, S3, IV fluids, vasopressors, diuretics/dialysis), learning cooperatively via a QMix network.
All models used same state representations, action discretization methods, and reward functions. Detailed feature processing information and model parameters are available in Appendix, Section Model Parameters. All codes for data processing and model training are included in Supplementary materials and will publicly available upon acceptance.
Results and Discussion
We present and discuss our experimental results in the form of answers to four key research questions:
RQ1: Does the proposed solution effectively learns a superior treatment policy?
We evaluate the performance of our learned policies using various off-policy quantitative and qualitative metrics (see Table 1) as follows:
Table 1: Comparison of off-policy evaluation metrics
Model ESS V Mortality (%)
Clinician 89115 18.98 16.27
D3QN-S 12 -1.21 18.69 ± 0.45
SoftAC-S 975 -0.19 18.52 ± 0.57
D3QN-O 99 0.37 14.47 ± 0.34
D3QN-T 373 19.42 14.02 ± 0.41
QMix-O 41 13.80 13.84 ± 0.15
QMix-T 602 25.83 11.29 ± 0.28
Proposed 7233 30.04 8.81 ± 0.24
1) Average Returns: The performance of learned policies is quantitatively evaluated using their estimated average returns. We use $V^{CWPDIS}$ ($V$), the average return from common sub-trajectories in both learned and clinician policies (Thomas 2015). Effective sample size ($ESS$) is the number of common sub-trajectories, capped at the total number in the clinician-policy. $ESS$ measures the confidence in the corresponding $V$ value. An effective policy should have a higher $V$ with a significant $ESS$. Evaluations showcase that all single agent baselines (D3QN-S and SoftAC-S) fail, resulting in negative average returns. This indicates the single agent systems’ inability to learn effective policies in this complex setting. The best performing baseline, QMix-T, had the highest ESS among the baselines and a larger $V$ than the clinician. Thus, while QMix-T’s policy does not fully align with the clinician’s, it can effectively select actions that result in larger immediate returns. Our proposed model considerably outperformed all baselines and the clinician policy, obtaining the highest $V$, supported by a significant $ESS$. This suggests that the proposed model learnt a superior policy which acts in accordance to high-return clinician actions while deviating from low return ones.
2) Mortality Rate: In line with literature, we estimate mortality rates using the clinician policy’s relationship between mortality and expected returns. Specifically, we categorize expected returns from patient trajectories into bins and calculate the mortality rate for each bin. The resulting relationship between the expected returns and mortality is used to estimate mortality rates for learned policies based on their expected returns. The single-agent baselines (D3QN-S and SoftAC-S) both showed an approximate 14% increase in mean mortality compared to the Clinician, indicating a failure to learn an improved policy. In contrast, cooperative multi-agent baselines (QMix-O and QMix-T) demonstrated a decrease of 14.9% and 30.6% in mortality. The non-cooperative multi-agent baselines (D3QN-O and D3QN-T) outperformed single agent baselines, but were less effective than the cooperative multi-agent baselines, highlighting the importance of collaboration among agents to obtain superior policies. Proposed model showcased the highest decrease in mean mortality by 45.9%, highlighting its ability to learn a policy that could considerably improve patient survival compared to both state-of-the-art models and clinician policy.
(a) D3QN-S
(b) SoftAC-S
@@
Figure 2: Mortality vs. expected return for all models. The shaded area represents standard errors.
3) Mortality vs. Expected Return: We further evaluate the efficacy of the learned policies by analyzing the correlation between mortality and expected returns (see Figure 2). An effective policy should display a strong negative correlation, where higher expected returns translate to lower mortality, and vice versa. This indicates that the policy learns to associate actions leading to lower returns with higher mortality, and vice versa. The multi-agent baselines (QMix-O and QMix-T) display steeper negative curves than the single-agent baselines (D3QN-S and SoftAC-S). However, these baselines either failed to maintain the negative correlation consistently throughout the plot with some positive correlations in some parts, or did not exhibit a steeper negative correlation. In contrast, the proposed model showed the steepest negative correlation with a consistent negative relationship, indicating the highest ability to enhance patient survival.
@@
Figure 3: Dosage differences (x-axis) versus mortality (y-axis) for the top single-agent baseline SoftAC (top), multi-agent baseline QMix-T (middle), and our proposed model (bottom), with colors blue, red, and green denoting different organ systems.
@@
4) Mortality vs. Difference in Recommended Dosage: For each intervention, we analyzed the relationship between mortality rates, and the recommended dosage differences between clinician-administered and learned policies. We estimated this relationship by categorizing quantile-level dosage differences into bins and computing mortality rates of each bin. An effective policy should align with clinician dosages that resulted in low mortality (x-axis=0), and increasingly deviate from those associate with increasing mortality, ideally forming a V-shaped curve centered at 0.
@@
We showcase the performance of our model against the best-performing single-agent (SoftAC-S) and multi-agent (QMix-T) models (see Figure 3). See Appendix, Section Experimental Results for comparison across all baseline models. The single-agent model failed to achieve the desired V-shape for any treatment, likely due to the large and complex action space. The multi-agent baseline shows comparatively better performance, demonstrating the advantage of cooperative agents operating under factored action spaces. Our model closely approximated the desired V-shape across treatments, except for S3 due to minimal sample sizes (see Appendix, Section RL Action for sample sizes).
Figure 4: Correlation matrices of 7 treatment types within clinician and proposed policies across low, medium, and high SOFA severity levels, categorized by (a) survived and (b) deceased outcomes.
RQ2: Do individual agents learn effective local policies?
In addition to evaluating the overall effectiveness of the learned hierarchical policy, we evaluated the local policies of individual agents within the proposed hierarchy, each operating within their factored state and action spaces. We analyzed mortality rates against clinician and agent recommended dosage differences (see Appendix, Section Experimental Results). All individual agents’ policies closely followed the desired ’V’ shape, indicating effective local policies across the hierarchy. This increases our confidence in the combined global policy and enhances the reliability of our proposed hierarchical task decomposition solution.

### v2

@@ 2.2.2. Actions
In addition to the agent options discussed above for higher-level agents, this section provides detailed information on the actions of leaf-level agents. According to our hierarchical task decomposition approach, each agent operates within a factored action space. This helps reduce the learning burden on each agent and prevents the effects of potentially low-quality samples across subspaces. Based on clinical expertise and data availability, we selected six treatments. Two for each the Cardiovascular system: IV fluids (IV) and vasopressors (Vaso), Neuro system: anesthetics (S1) and analgesics (S2), and Renal system: diuretics and dialysis. These organ systems and their respective treatments were chosen by a physician according to clinical priority in intensive care patients. Each treatment plays a unique and critical function within its biological system and requires careful control to ensure efficacy and safety. Our framework is scalable to any number of organ systems and treatments. The proposed hierarchy is generic which allows integration of agents operating on continuous and discrete action spaces. However, we tested the approach using discrete action spaces. Dosages for each treatment were discretized into five levels (no-action + 4 quantiles), except dialysis, which is a binary action (active or inactive). Proposed hierarchy supports mixing of any treatment within or across systems, aligning with common clinical practices derived from our data analysis and clinical consultations. For example, we enforced exclusive use of one renal treatment at a time—either diuretics or dialysis, and allowed all other treatment combinations. This showcases the hierarchy’s adaptability to incorporate such domain specific constrains. Accordingly, our solution flexibly handles organ systems with different degrees of complexity and configurations of actions, enabling the effectiveness of the solution in real, complex clinical settings.
2.2.3. Reward Function
We used a hybrid reward system, combining a mortality-based terminal reward—positive (+R+R) for survival, and negative (R-R) for death—with clinically guided intermittent rewards to adjust rewards based on immediate health outcomes. Specifically, we define intermediate rewards based on changes in the Sequential Organ Failure Assessment (SOFA) score and blood lactate levels—two validated surrogates of physiological deterioration. The reward penalizes both persistently high SOFA scores and increases in SOFA or lactate between successive time steps, while assigning positive reward for improvements. The overall form is:
(1) $R(s_{t},s_{t+1})=\begin{cases}+R,&\text{if }s_{t+1}\in\mathcal{S}_{T}\cap\mathcal{S}_{sur}\\ -R,&\text{if }s_{t+1}\in\mathcal{S}_{T}\cap\mathcal{S}_{dec}\\ R_{im}(s_{t},s_{t+1}),&\text{if }s_{t+1}\notin\mathcal{S}_{T}\end{cases}$
(2) $\begin{split}R_{im}(s_{t},s_{t+1})=&\;C_{0}\big{(}s_{t+1}^{SOFA}=s_{t}^{SOFA}\;\land\;s_{t+1}^{SOFA}>0\big{)}\\ &+C_{1}\big{(}s_{t+1}^{SOFA}-s_{t}^{SOFA}\big{)}\\ &+C_{2}\tanh\big{(}s_{t+1}^{lactate}-s_{t}^{lactate}\big{)}\end{split}$
where $C_{0}=-0.025$, $C_{1}=-0.125$, and $C_{2}=-2$. The $\tanh$ term ensures that large lactate swings do not disproportionately affect learning.
The chosen approach addresses the challenges of credit assignment and reward sparsity in long sequences. They ensure that immediate health improvements are reflected in decision-making while supporting adaptable, cross-organ policy learning applicable across diverse healthcare settings. Our reward design follows established structures in healthcare RL research; however our framework itself is generalizable as it is not tied to an specific reward function. Where terminal rewards suffice, our framework can operate without intermittent rewards.
2.3. Q Learning
@@
Traditional RL frameworks, using Markov Decision Processes (MDPs), often assume actions are executed instantaneously and have uniform durations. This simplifies modeling, but fails to capture real-world complexities with actions spanning multiple steps and varying duration.
@@
where $\gamma$ is a discount factor and $\epsilon^{\pi}(o,s,t)$ is the execution of option $o$ in state $s$ from time $t$ to its termination. Unlike the standard options framework, our solution contains independent and cooperative agents controlling options (see Section 2.1). Thus, root level Q-values are updated as follows:
(3) $Q(s,o)\leftarrow Q(s,o)+\alpha\left[Q_{tgt}(s,o)-Q(s,o)\right]$
@@
[…]
where $r$ is the immediate reward for selecting no-action $A_{Rt}^{0}$, $A_{Neu}$, $A_{Car}$, $A_{Ren}$ and $A_{Rt}^{OMix}$ are the action spaces for Neuro, Cardio, Renal agents, and their combined action space $A_{Neu}\times A_{Car}\times A_{Ren}$. Value functions $Q_{Neu}(s,a)$, $Q_{Car}(s,a)$, and $Q_{Ren}(s,a)$ for corresponding agents are updated using Temporal Difference (TD) learning as follows:
(4) $Q_{Neu}(s,a)\leftarrow$ $Q_{Neu}(s,a)+\alpha\left[y^{Neu}-Q_{Neu}(s,a)\right]$
@@
$M_{OMix}$ is trained using the QMix architecture (Rashid et al., 2020), combining individual value functions of its sub-agents ($M_{Omix}^{Neu}$$M_{Omix}^{Car}$, and $M_{Omix}^{Ren}$) into a unified value function ($Q_{OMix}$). QMix supports cooperative training and consistent decision-making across agents by enforcing a monotonicity constraint. The constraint keeps the weights of the mixing network non-negative, ensuring that the argmaxargmax operation on $Q_{OMix}$ is consistent with those from the sub-agents. Hence, $Q_{OMix}$ is represented as follows:
(5) $\mathrm{arg\,max}_{a\in A_{OMix}}Q_{OMix}(s,a)=$
$\left(\begin{aligned} &\mathrm{arg\,max}_{a_{OMix}^{Neu}\in A_{OMix}^{Neu}}Q_{OMix}^{Neu}(s_{OMix}^{Neu},a_{OMix}^{Neu})\\ &\mathrm{arg\,max}_{a_{OMix}^{Car}\in A_{OMix}^{Car}}Q_{OMix}^{Car}(s_{OMix}^{Car},a_{OMix}^{Car})\\ &\mathrm{arg\,max}_{a_{OMix}^{Ren}\in A_{OMix}^{Ren}}Q_{OMix}^{Ren}(s_{OMix}^{Ren},a_{OMix}^{Ren})\end{aligned}\right)$
@@
2.3.1. Training Process:
We used a two phase approach to train the proposed hierarchical model (see Algorithm 1 in Appendix). In phase one, the root agent ($M_{Rt}$) is trained first. Resulting latent embeddings $E^{Rt}$ are then fine-tuned during the subsequent training of organ-specific agents ($M_{Neu}$, $M_{Car}$, and $M_{Ren}$). In the process, targeted embeddings ($E^{Neu}$, $E^{Car}$ and $E^{Ren}$) are learned and formulate Targeted State Representations. These representations are used to train the corresponding lower-level sub-agents. Independent agents directly learn $Q$ values from received rewards (see Equation 4), whereas cooperative agents $M_{Omix}^{Neu}$$M_{Omix}^{Car}$, $M_{Omix}^{Ren}$ are trained using shared rewards (Equation 5). In phase two, we integrate the full hierarchy by using the trained organ-specific agents from phase one. $Q(s,a)$ values from these agents are used to retrain $M_{Rt}$ (see Equation 3), and no other agent is retrained. After training, at each timestep, $M_{Rt}$ evaluates the patient state $s_{t}^{Rt}$ and select either no-action, or invokes lower level agents for final treatment and dosage recommendations.
3. Experiments
3.1. Datasets
We collected data on 30,440 patients under Sepsis-3 criteria from the popular MIMIC-IV database (Johnson et al., 2020). The data spans from 24 hours pre-diagnosis to 48 hours post-diagnosis forming a maximum 72 hour window per patient, barring death or ICU discharge. We defined a continuous state-space which included 44 physiological measurements including vital signs, laboratory test results, severity scores and demographics. State and actions data were aggregated into four-hourly windows to generate uniform patient data sequences. The patient trajectories were randomly split into 75% training and 25% testing sets. Additionally, we collected data for 4,390 sepsis patients from the AmsterdamUMC dataset (Thoral et al., 2021) to conduct an external validation of our trained models using the same set of state and action variables. High level statistics of the cohorts are listed in Table 1. See Appendix, Section RL Actions and RL States for more information on the state and action spaces.
Table 1. Cohort characteristics for MIMIC-IV and AmsterdamUMC
Dataset Cohort % Female Age (Mean) ICU LOS (Mean) Population
MIMIC-IV Overall 42.11 66.14 5d 4h 30,440
Non-Survivors 44.00 69.77 6d 12h 4,948
Survivors 41.75 65.44 4d 22h 25,492
AmsterdamUMC Overall 38.82 61.71 7d 8h 4,390
Non-Survivors 39.49 66.64 9d 9h 961
Survivors 38.63 60.33 6d 18h 3,429
3.2. Baselines
We evaluated our model performance against multiple baselines, including single- (D3QN-S and SoftAC-S) and multi-agent systems under independent (D3QN-O) and cooperative (QMix-O and QMix-T) learning approaches:
@@ 3.2. Baselines
  • Single D3QN (D3QN-S) (Raghu et al., 2017): Single-agent D3QN predicting all treatment and dosage combinations in a flattened action space. We used the original implementation with tuned hyperparameters.
  • Single SoftAC (SoftAC-S) (Haarnoja et al., 2018): A single-agent Soft Actor-Critic predicting all action combinations in a flattened action space. We used the original implementation and tuned hyperparameters.
  • Organ-specific D3QN (D3QN-O): Three independent D3QN agents for Neuro, Cardio and Renal systems. Models are trained using all available samples, including those treated exclusively for the target organ and those mixed with other organ treatments. Averaged quantitative metrics are reported.
  • Treatment-specific D3QN (D3QN-T): Five independent D3QN agents for S1, S2, IV, vaso, and diuretics/dialysis. Models are trained using all available samples, including those using single-treatments and mixed-treatments. Averaged quantitative metrics are reported.
@@
  • Treatment-coordinated QMix (QMix-T): Uses five cooperative treatment-level agents (S1, S2, IV fluids, vasopressors, diuretics/dialysis), learning cooperatively via a QMix network.
All models used the same state representations, action discretization methods, and reward functions. All codes for data processing and model training are included in the supplementary materials and will publicly available upon acceptance.
3.3. Results and Discussion
We present and discuss our experimental results as answers to four key research questions as follows.
Table 2. Policy Evaluation on MIMIC-IV and AmsterdamUMC
Policy MIMIC-IV External Validation
V Mortality (%) V Mortality (%)
Clinician 9.08 16.27 7.74 21.89
SoftAC-S -3.26 20.52 $\pm$ 0.68 -6.22 28.59 $\pm$ 0.85
D3QN-S -1.29 19.79 $\pm$ 0.77 -6.87 29.13 $\pm$ 0.92
D3QN-O 12.74 14.57 $\pm$ 0.30 8.01 20.57 $\pm$ 0.50
D3QN-T 13.05 13.62 $\pm$ 0.43 10.24 19.91 $\pm$ 0.45
QMIX-O 15.45 13.29 $\pm$ 0.37 8.70 20.12 $\pm$ 0.55
QMIX-T 17.16 12.74 $\pm$ 0.30 11.07 18.56 $\pm$ 0.46
Proposed 21.73 9.95 $\pm$ 0.33 12.95 17.34 $\pm$ 0.43
(a) SoftAC-S
(b) D3QN-S
@@ V Mortality (%) V Mortality (%)
Figure 2. Mortality vs. expected return for all models. The shaded area represents standard errors.
3.3.1. RQ1: Does the proposed solution effectively learns a superior treatment policy?
We evaluate the performance of learned policies using off-policy quantitative and qualitative metrics (see Table 2). They do not require a simulator, but adopt a well-established approach using offline learning on retrospective data. These metrics, widely used in healthcare RL enable safe, effective learning from clinical outcomes without real-time exploration risks (Raghu et al., 2017; Komorowski et al., 2018).
@@ 3.3.1. RQ1: Does the proposed solution effectively learns a superior treatment policy?
Figure 3. Dosage differences (x-axis) versus mortality (y-axis) for the top single-agent baseline SoftAC, multi-agent baseline QMix-T, proposed model on the testing set, and followed lastly by the proposed model on the external validation set, with colors blue, red, and green denoting different organ systems.
1) Average Returns: We use Consistent Weighted Per-Decision Importance Sampling (CWPDIS), a standard OPE method to estimate expected returns from each policy $V^{CWPDIS}$ ($V$). (Thomas, 2015). A higher $V$ corresponds to a more effective policy. Both single agent baselines (D3QN-S and SoftAC-S) result in negative average returns. This indicates the single agent systems’ inability to learn effective policies in this complex setting. Meanwhile, the multi-agent model QMix-T was the best performing baseline according to OPE returns. This suggests that for our particular multi-organ problem, MARL approaches with some form of cross-agent communication can be beneficial for policy learning. Overall, our model considerably outperformed all baselines and the clinician policy, obtaining the highest estimated returns ($V=21.73$). This suggests that the model learnt a superior policy which acts in accordance to high-return clinician actions while deviating from low return ones. Results of the external validation show consistent, albeit more modest results. The smaller improvements in the OPE returns on the AmsterdamUMC dataset may be explained by differences in patient populations (U.S vs European cohorts) and the differences in clinical setting/protocols, among other reasons.
2) Mortality Rate: In line with literature, we estimate mortality rates using the clinician policy’s relationship between mortality and expected returns. We do this by categorizing expected returns from patient trajectories into bins and calculate the mortality rate for each bin. The resulting relationship between the expected returns and mortality is used to estimate mortality rates for learned policies based on their expected returns. The single-agent baselines (SoftAC-S and D3QN-S) showed approximate 26% and 22% increases in mean mortality rate compared to the clinician respectively, indicating a failure to learn an improved policy. In contrast, cooperative multi-agent baselines (QMix-O and QMix-T) showed a decrease of 12.2% and 21.7% in mortality rate. The non-cooperative multi-agent baselines (D3QN-O and D3QN-T) outperformed single agent baselines, but were less effective than the cooperative multi-agent baselines, highlighting the importance of collaboration among agents to obtain superior policies. Our model had the highest decrease in mean mortality by 38.9%, highlighting its ability to learn a policy that could considerably improve patient survival compared to state-of-the-art models and clinician policy. While we see a relatively smaller improvement in mortality rate in our external validation (20.9% decrease), our proposed model still achieves the lowest estimated mortality rate among all baselines, supporting the performance advantages of our approach.
3) Mortality vs. Expected Return: We further evaluate the efficacy of the learned policies by analyzing the correlation between mortality and expected returns (see Figure 2). An effective policy should display a strong negative correlation, where higher expected returns translate to lower mortality, and low expected returns to higher mortality. This relationship reflects the policy’s ability to discern optimal actions from sub-optimal actions. The multi-agent baselines (QMix-O and QMix-T) display steeper negative curves than the single-agent baselines (D3QN-S and SoftAC-S). However, overall we see that none of the baselines consistently maintained the negative correlation throughout the plot—with plateauing and positive correlations at some points in the plot. In contrast, the proposed model showed the steepest negative correlation with a consistent negative relationship, indicating the highest ability to improve patient survival. In our external validation, we see the same consistent negative relationship, indicating a good degree of generalizability.
@@
4) Mortality vs. Difference in Recommended Dosage: For each intervention, we analyzed the relationship between mortality rates, and the recommended dosage differences between clinician-administered and learned policies. We estimated this relationship by categorizing quantile-level dosage differences into bins and computing mortality rates of each bin. An effective policy should align with clinician dosages that resulted in low mortality (x-axis=0), and increasingly deviate from those associate with increasing mortality, ideally forming a ’U’ or ’V’-shaped curve centered at 0.
@@
We showcase the performance of our model against the best-performing single-agent (D3QN-S) and multi-agent (QMix-T) models (see Figure 3). See Appendix, Section Experimental Results for comparison across all baseline models. The single-agent model failed to achieve the desired V-shape for any treatment, indicating a failure to properly learn from the large action space. The multi-agent baseline shows comparatively better performance, again demonstrating the advantage of cooperative agents operating under factored action spaces. On the other hand, our proposed model most closely approximated the desired V-shaped relationship across treatments; these qualitative results were also adequately replicated in our external validation, highlighting decent generalizability of our trained model.
Figure 4. Mortality versus dosage difference (Clinician – Model) for individual agents across Neuro (blue), Cardio (red), and Renal (green) systems. Rows represent: individual treatments (top), intra-organ mixtures (middle), and inter-organ coordination (bottom). Renal agents do not include intra-organ mixtures due to clinical exclusivity between diuretics and dialysis.
3.3.2. RQ2: Do individual agents learn effective local policies?
To assess if the agents operating within localized state-action subspaces learn effective treatment policies, We analyzed the relationship between mortality rates and discrepancies between the clinician and each agent recommended dosages (see Figure 4). Across all three organ systems, we observe that most agents showcase the desired V shape. For example, the Vasopressors and S1 agents show steep increases in mortality when deviating from clinician dosages, showing the sensitivity of the interventions. Proposed model tend to match effective dosing and adaptively diverge in high-risk scenarios, which supports is clinical relevance. Importantly, the renal sub-agents only appear in individual and cross-organ coordination rows as they do not contribute to intra-organ mixtures. This was intentionally modeled as renal treatments were found mutually exclusive in clinical practice, where diuretics and dialysis are found to be applied at different severity levels and rarely administered together. Enforcing exclusivity improves training stability and better reflects real-world treatment logic and flexibility in the proposed framework.